You react 10.0 g of nitrogen gas with hydrogen

gas according to the following reaction.

N 2 (g) 3 H2 (g) n 2N H3 (g)


a. What mass of hydrogen is required to

completely react with 10.0 g sample of

nitrogen gas?

b. What mass of ammonia is produced from

10.0 g of nitrogen gas and sufficient

hydrogen gas?


1
Expert's answer
2021-03-10T06:52:04-0500

(a) moles of N2 "=\dfrac{mass}{MM}= \dfrac{10.0g}{28.02g/mol}=0.3569mol"


Moles of H2 needed will then be;


Moles ratio of N2:H2 = 1:3


Moles of H2 "=\dfrac{3}{1}x0.3569mol = 1.0707mol"


Mass of H2 = moles x MM = 1.0707mol x 2.016g/mol

= 2.158g


(b) Moles of N2 = 0.3569mol

Mole ratio of N2 : NH3 = 1:2


Moles of NH3 "=\dfrac{2}{1}x0.3569mol = 0.7138mol"


Mass of NH3= moles x MM

= 0.7138mol x 17.031g/mol

=12.134g